# Exploring the Privacy-Energy Consumption Tradeoff for Split Federated Learning

### v3

@@ III-A An Illustrative Example
As shown in Fig. 4, increasing the depth of the cut layer (i.e., increased complexity of the client-side model) leads to higher energy consumption ($E$), causing both $E_{\text{main}}$ and $E_{\text{fed}}$ to rise due to the increased computing and networking energy usage. Concurrently, RS consistently decreases with deeper cut layers. This phenomenon implies that the depth of the cut layer can be directly correlated with the complexity of the client-side model, potentially making attackers to reconstruct the image from the smashed data more difficult. Nevertheless, due to the lack of analytic understanding of RS, a data-driven approach can be applied to model RS through offline training using empirical measurements. In this context, we adopt regression-based modeling, which is one of the most widely used approaches, including mobile CPU property models (e.g., CPU power and temperature variation modeling). Consequently, we demonstrated that a convex model can effectively approximate the behavior of RS, yielding a Root Mean Square Error (RMSE) of only $0.0028$. As a result, RS can be estimated as $\operatorname{RS}(\alpha)$ = $0.76750.3597\alpha^{2}-0.7004\alpha+$ 0.7675. It should be noted that the estimated $RS(\alpha)$ may differ across applications; therefore, it should be pre-calculated in an offline manner prior to making any adjustments. Consequently, we can optimize the selection of the cut layer, denoted by $\alpha$, to minimize privacy leakage from reconstruction attacks, as measured by $RS(\alpha)$, while ensuring that energy consumption ($E$) remains within an acceptable energy budget ($E_{\text{req}}$). Thus, the problem can be formulated as follows:
@@
III-B Other Types of Control Variables
  • Computation/Power/Radio resource management: The computation resource, transmission power, and radio resource (i.e., bandwidth allocation over clients) of each client participating in SFL plays a pivotal role in optimizing performance. However, due to the early stage of this field, there has been a lack of rigorous analysis for SFL. Fundamentally, those factors are known to affect overall latency, denoted as $T_{i}$ in our analysis, which in turn impacts energy consumption within the field of FL [10].
@@ III-B Other Types of Control Variables
As we reviewed, the performance of SFL, concerning both energy efficiency and convergence speed, etc., is noteworthy. The heterogeneity in clients’ computing and networking capabilities can impact the speed of convergence due to the issue of stragglers, and this diversity also leads to differences in energy consumption among clients. Correspondingly, achieving an optimal balance concerning these multiple variables often requires substantial complexity in pursuing a global optimum. This necessitates adopting a modular design strategy or employing various heuristic methods to achieve sub-optimal solutions.
IV Open Challenges
IV-A Deep Reinforcement Learning
@@
Considering the dynamic and heterogeneous nature of client networking and computing capabilities, deep reinforcement learning (DRL) shows promise in intelligently selecting the optimal cut layer. This aims to balance privacy, energy consumption, and convergence speed in the SFL, along with other essential control parameters, as discussed. Recent research by [15] illustrated how DRL is utilized to optimize cut layer selection for collaborative DL inference in multi-access computing (MEC) servers by efficiently offloading certain inference tasks from clients to the MEC server. The key challenges in applying DRL to address various SFL issues revolve around precisely defining the agent, environment, state, action, and reward. Additionally, selecting an appropriate DRL model and fine-tuning hyperparameters become pivotal for improving convergence speed, especially when dealing with complex and dynamic environments.
IV-B Privacy and Security Protection
@@
As in [9], when clients and the main server hold their data and labels, respectively, attackers could be malicious data owners or eavesdroppers trying to intercept gradients shared from the main server to clients. The similarity in cut layer gradients among data samples can disclose their connection to labels, known as a label inference attack. Therefore, to mitigate risks, label inference attacks must be addressed by carefully choosing the cut layer and implementing differential privacy (DP). Moreover, determining the cut layer is also critical for maintaining security and integrity, particularly concerning model poisoning and data integrity. Specifically, the cut layer selection significantly influences the system’s susceptibility to model poisoning attacks. For instance, extensive client-side training elevates the risk of client-side poisoning, complicating the defense against distributed attacks. Thus, sophisticated aggregation techniques are required to counter model poisoning risks effectively. An inappropriately chosen cut layer can also hinder data authenticity and integrity validation, posing challenges to maintaining the model’s training integrity. The selection of the cut layer, therefore, demands careful consideration, balancing between minimizing security risks and maximizing the efficiency of SFL.
IV-C Lightweight Design
@@
To enhance the efficiency of SFL, it is beneficial to consider the quantization approach. In the context of SFL, adopting a quantization approach can substantially reduce the size of both local model updates and intermediate data, which is useful particularly in resource-constrained clients. This reduction has several advantages, including reduced bandwidth usage, lower transmission energy, and reduced storage needs. Furthermore, quantization introduces inherent privacy guarantees. This is because reducing the precision of transmitted data inherently obfuscates the exact values, thereby adding an extra layer of protection for sensitive information, which is crucial in maintaining data privacy. Moreover, to mitigate the burden of training on resource-constrained clients with reduced energy consumption, the strategy of downsizing the resolution and the total volume of training data can be employed, considering different data distributions and acceptable accuracy levels. Additionally, selecting a lightweight DL model or employing model pruning techniques is adopted to facilitate a lightweight design for SFL.
V Conclusions
@@
This article has provided an overview of the SFL, mainly focusing on its communication efficiency and privacy issues. By studying the impact of the cut layer selection on both energy consumption and privacy, we have provided a concrete example of an efficient cut layer selection to minimize the risk of reconstruction attacks within the required energy budget. Finally, we have suggested various other adjustable factors and highlighted some promising research directions within this field.
@@
Mohamed Seif (mseif@princeton.edu) is currently a Post-Doctoral Research Associate with Princeton University, Princeton, NJ, USA. His current research interests include information theory, machine learning, and wireless communications.
Jungchan Cho (thinkai@gachon.ac.kr) is currently an Assistant Professor in the School of Computing at Gachon University, Seongnam, South Korea. From 2016 to 2019, he was a Senior Software Engineer at Samsung Electronics. His research interests include deep learning, computer vision, and machine learning.

### v4

@@ 3.1 An Illustrative Example
As shown in Fig. 4, increasing the depth of the cut layer (i.e., increased complexity of the client-side model) leads to higher energy consumption ($E$), causing both $E_{\text{main}}$ and $E_{\text{fed}}$ to rise due to the increased computing and networking energy usage. Concurrently, RS consistently decreased with deeper cut layers. This phenomenon implies that the depth of the cut layer can be directly correlated with the complexity of the client-side model, potentially making attackers to reconstruct the image from the smashed data more difficult. Nevertheless, owing to the lack of an analytical understanding of RS, a data-driven approach can be applied to model RS through offline training using empirical measurements. In this context, we adopted regression-based modeling, which is one of the most widely used approaches, and includes mobile CPU property models (e.g., CPU power and temperature variation modeling). Consequently, we demonstrate that a convex model can effectively approximate the behavior of RS, yielding a Root Mean Square Error (RMSE) of only $0.0028$. Consequently, RS can be estimated as $\operatorname{RS}(\alpha)$ = $0.76750.3597\alpha^{2}-0.7004\alpha+$ 0.7675. Note that the estimated $\operatorname{RS}(\alpha)$ may differ across applications; therefore, it should be pre-calculated offline before making any adjustments. Consequently, we can optimize the selection of the cut layer, denoted by $\alpha$, to minimize privacy leakage from reconstruction attacks, as measured by $RS(\alpha)$, while ensuring that energy consumption ($E$) remains within an acceptable energy budget ($E_{\text{req}}$). Therefore, the problem can be formulated as follows:
@@
3.2 Other Types of Control Variables
  • Computation/Power/Radio resource management: The computation resource, transmission power, and radio resource (i.e., bandwidth allocation over clients) of each client participating in SFL play a pivotal role in optimizing performance. However, due to the early stage of this field, there has been a lack of rigorous analysis for SFL. Fundamentally, those factors are known to affect overall latency, denoted as $T_{i}$ in our analysis, which in turn impacts energy consumption within the field of FL [13].
@@ 3.2 Other Types of Control Variables
As reviewed, the performance of SFL in terms of both energy efficiency and convergence speed is noteworthy. Heterogeneity in clients’ computing and networking capabilities can affect the speed of convergence owing to the issue of stragglers, and this diversity also leads to differences in energy consumption among clients. Correspondingly, achieving an optimal balance among these multiple variables often requires substantial complexity when pursuing a global optimum. This necessitates the adoption of a modular design strategy or the employment of various heuristic methods to achieve suboptimal solutions.
4 Open Challenges
4.1 Deep Reinforcement Learning
@@
Considering the dynamic and heterogeneous nature of client networking and computing capabilities, deep reinforcement learning (DRL) shows promise in intelligently selecting an optimal cut layer. This aims to balance privacy, energy consumption, and convergence speed in the SFL, along with other essential control parameters. A recent study by [15] illustrated how DRL is utilized to optimize cut layer selection for collaborative DL inference on multi-access computing (MEC) servers by efficiently offloading certain inference tasks from clients to the MEC server. The key challenges in applying DRL to address various SFL issues involve precisely defining the agents, environments, states, actions, and rewards. In addition, selecting an appropriate DRL model and fine-tuning hyperparameters become pivotal for improving convergence speed, particularly when dealing with complex and dynamic environments.
4.2 Privacy and Security Protection
@@
As in [12], when clients and the main server hold their data and labels, respectively, attackers can be malicious data owners or eavesdroppers trying to intercept the gradients shared between the main server and clients. The similarity in cut layer gradients among data samples can reveal their connection to labels, which is known as a label-inference attack. Therefore, to mitigate risks, label-inference attacks must be addressed by carefully choosing the cut layer and implementing differential privacy (DP). Moreover, determining the cut layer is critical for maintaining security and integrity, particularly for model poisoning and data integrity. Specifically, the cut layer selection significantly influences a system’s susceptibility to model-poisoning attacks. For instance, extensive client-side training increases the risk of client-side poisoning, complicating the defense against distributed attacks. Thus, sophisticated aggregation techniques are required to effectively counter poisoning risks. An inappropriately selected cut layer can also hinder data authenticity and integrity validation, posing challenges to maintaining the model’s training integrity. Therefore, the selection of the cut layer requires careful consideration, balancing minimizing security risks and maximizing the efficiency of SFL.
4.3 Lightweight Design
@@
To enhance the efficiency of SFL, it is beneficial to consider a quantization approach. In the context of SFL, adopting a quantization approach can substantially reduce the size of both local model updates and intermediate data, which is particularly useful for resource-constrained clients. This reduction has several advantages, including reduced bandwidth usage, lower transmission energy, and reduced storage requirements. Furthermore, quantization introduces an inherent privacy guarantee. This is because reducing the precision of the transmitted data inherently obfuscates exact values, thereby adding an extra layer of protection to sensitive information, which is crucial for maintaining data privacy. Moreover, to mitigate the burden of training on resource-constrained clients with reduced energy consumption, a strategy of downsizing the resolution and total volume of the training data can be employed, considering different data distributions and acceptable accuracy levels. In addition, the selection of a lightweight DL model or employing model-pruning techniques was adopted to facilitate a lightweight design for SFL.
5 Conclusions
@@
This article provides an overview of SFL, focusing primarily on communication efficiency and privacy issues. By studying the impact of cut layer selection on both energy consumption and privacy, we provided a concrete example of efficient cut layer selection to minimize the risk of reconstruction attacks within the required energy budget. Finally, we suggest various other adjustable factors and highlight the promising research directions in this field.
Acknowledgments
This work was supported by the U.S National Science Foundation under Grant ECCS-2335876.
@@ Acknowledgments
Mohamed Seif (mseif@princeton.edu) is currently a Post-Doctoral Research Associate at Princeton University. His current research interests include information theory, machine learning, and wireless communications.
Jungchan Cho (thinkai@gachon.ac.kr) is currently an Associate Professor in the School of Computing at Gachon University. From 2016 to 2019, he was a Senior Software Engineer at Samsung Electronics. His research interests include deep learning, computer vision, and machine learning.